OpenStudy (anonymous):

will medal Solve 2x2 − 6x = −8. negative 3 plus or minus 5i all over 2 the quantity of 3 plus or minus 2i square root 30 all over 2 the quantity of 3 plus or minus i square root 30 all over 2 the quantity of 3 plus or minus i square root 7 all over 2

OpenStudy (johnweldon1993):

Lets complete the square and find out We need to x^2 term to have a coefficient of 1...so lets divide everything by 2 \[\large x^2 - 3x = -4\] Now...lets take half the coefficient of the 'x' term....so 3/2 and square that....so 9/4...and add this to both sides of our equation \[\large x^2 - 3x + \frac{9}{4} = -4 + \frac{9}{4}\] Simplify the right hand side...and turn the left hand side into a square \[\large (x - \frac{3}{2})^2 = -\frac{7}{4}\] Take the square root of both sides \[\large x - \frac{3}{2} = \sqrt{-\frac{7}{4}}\] Add \(\large \frac{3}{2}\) to both sides \[\large x = \frac{3}{2} \pm \sqrt{-\frac{7}{4}}\] Which can then be written as \[\large x = \frac{3}{2} \pm \frac{\sqrt{-7}}{2}\] Now how else can we simplify that?

OpenStudy (anonymous):

by adding them ?